OpenStudy (anonymous):

Solve the problem by stating the equation you would use to solve and then solving it. Be sure to show all of your work. In a local election, 25,800 people voted. This was an increase of 5% over the last election. How many people voted in the last election? Round to the nearest whole person.

OpenStudy (anonymous):

let number voted in last election = x then x + 0.05x = 25,800 1.05x = 25800 x = 25800 / 1.05 = 24,571 to nearest person